if one of the root of the equation 8x square-6x-a-3=0  is the square of the other , then the values of a are   (a)- 4,-24  (b)-  4,24  (c)-   -4,-24  (d)-    -4,24

Grade:12th pass

1 Answers

Charchit Tailong
8 years ago
root1=p
root2=q

pq=(-a-3)/8
p+q=6/8
p^2=q